Seafoam:

Complete the square to rewrite the following equation. Identify the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 − 6y = −4

Seafoam:

@AZ

Florisalreadytaken:

as long as you are asking this question, you at least know something about the equation of a circle right?

Seafoam:

Afraid not

snowflake0531:

We can separate this to x^2 +4x + ___ y^2 - 6y + ___ The 2 __s can later be added to the other side We need this as (_ +/- _)^2 so what to the power of two would bring us x^2 +4x + something, and what would bring us to y^2 - 6y + something

snowflake0531:

Hint: x^2 + 4x + __ = (x+2)^2 y^2 -6y + ___ = (x-3)^2 So fill in the gaps, what are the values

snowflake0531:

Because I already gave you the other side (x+2)^2 and (x-3)^2 as hints you can factor it out and find the blanks by just doing (a+b)^2 = a^2 +2ab+b^2 (a-b)^2 = a^2 -2ab + b^2

snowflake0531:

And since I only want the last blank, from (x+2)^2 and (x-3)^2 just do 2^2 and -3^2

Seafoam:

2 squared would be 4 and -3 squared would be 9

snowflake0531:

Yes so that means we have x^2 + 4x+ 4 + y^2 -6x +9 and because we added the 4 and 9 to the right side, we have to add it back to hte right (x^2 + 4x+4) + (y^2 -6x +9) = -4 + 4+9 Can you simplify the right, and factor the left?

Seafoam:

3, (-2,3)

snowflake0531:

.... that's the answer yes, but you know that because I told you e.e

snowflake0531:

We have the left, and then we see the right 4-4+9, which is 9 SO, (x^2 + 4x+4) + (y^2 -6x +9) = 9

snowflake0531:

Since the right side is looking at radius, 9 is 3^2, which means that the radius is 3, since in (x-h)^2 + (y-k)^2 = r^2, r is radius, (h,k) is the center

Seafoam:

So the diameter is 9?

snowflake0531:

So, x^2 + 4x+4 is (x+2)^2 and (y^2 -6x+9) is (y-3)^2 so we have (x+2)^2 + (y-3)^2 = 3^2

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Seafoam So the diameter is 9? \(\color{#0cbb34}{\text{End of Quote}}\) ... diameter is 2 times radius, and radius is 3 e.e

Seafoam:

Oh shoot

snowflake0531:

Identify the center and radius of the circle. So we have radius is 3

snowflake0531:

Then since we have (x+2)^2 + (y-3)^2 this means that the coordinate is (-2,3), because you have to flip the signs

Seafoam:

And you just have to remember to flip the signs here I suppose?

snowflake0531:

Yes, for example if it was (x-2) and then (y+4) The coordinate would be (2,-4)

snowflake0531:

Do you... kinda..g et how to do it? lol

Seafoam:

@snowflake0531

snowflake0531:

Factoring: Assuming you know how to do FOIL, we know that (a+b)(c+d) = ac+ad+bc+bd Remember bd For example, we have x^2 + 10x + 16, we want to factor this Looking at x^2 we know that it will have to be (x+ _ )(x+ _ ) We can then look at c, (in ax^2 + bx +c) and see it as 16 We see b is a positive number, so we don't need to worry about signs, but remember that is as 10x, which means that the two numbers that we need to multiply to get 16 (bd) has to add up to 10 So we can list it out 1,16 2,8 3,6 4,4 Only one pair, 2 and 8, can add up to 10x So we know that is these two factors of 16 So then we have (x+2)(x+8) You can also recheck using FOIL `So, when you do those that have b and c both positive first look at the coefficient of x^2, usually when you practice factoring, it is 1, which makes it easier, as it is just (x+ _ )(x+ _ ) Then we look at C, know that 2 factors of it has to have a product of that, but also add up to the b value` Then, we think, what if the b value is negative, but c value is positive Example: x^2 - 8x + 16 We know that two negatives make a positive, so the only thing that matters, is change the two plus' into minus, from (x+ _ )(x+ _ ) into (x- _ )(x- _) So, by using this, the negative in front of b doesn't matter anymore We list the factors of 16, and see which ones add up to 8 1, 16 2, 8 3, 6 4, 4 We can see that the pair is 4 and 4 So plugging those in, it's (x-4)(x-4),, you can also write it as (x-4)^2 if you want `So, when you do these cases, change it to the form (x- _ )(x- _ ) And ignore the negative sign in front of the value b` The other kinds: for example, you could have x^2 +x -12 We don't know the signs yet but can first set it out (x +/- _ )(x +/- _ ) We see -12, but a positive value for b, (1x) This means that with the two factors of -12, with one positive and one negative, the larger one must be positive to make b positive So, we list them out 1,12 2,6 3,4 Not looking at the signs first, we can immediately cross out the first two pairs of factors, because no matter which is negative which is positive, it won't be x For this time, I'll write them out, with the larger positive -1,12 -> 11 --> 11x -2,6 -> 4 ----> 4x So, we know that it is 3 and 4, and with what I said earlier, 4 is positive, 3 is negative, because 4 plus -3 is x, which gets what we want So we have x^2 + x -12 ---> (x-3)(x+4) You can check using foil So in general When you do these 1. Check positive and negative for b and c (in ax^2 + bx +c ) 2. Check the C value, write down the factors 3. Look at b value, and see which pair adds up to the b value

snowflake0531:

e.e you shoulda told me that you didn't know FOIL FOIL = forward outside inside last Like in (a+b)(c+d) forward is a times c, ac, outside is a times d, ad, inside is b times c, bc, and last is b times d, bd so (a+b)(c+d) = ab+ad+bc+bd

Seafoam:

@snowflake0531

snowflake0531:

Okay so start with x^2 + 4x + y^2 − 6y = −4 Separate to (x^2 + 4x) + (y^2-6y) = -4 Want to make them trinomials to factor, something squared (x^2+4x+4) + (y^2 - 6y+9) = -4 + 4+9 Since I have to add to both sides^^^ Then we can change the trinomials, factor, and simplify the right (x+2)^2 + (y-3)^2 = 9 So then it's (x-h)^2 + (y-k)^2 =r^2 (x+2)^2 + (y-3)^2 = 3^2 where (h,k) is the center, r is radius So then we have (-2,3) is center, and 3 is radius

Seafoam:

Alright so can I answer the question like this x^2 + 4x + y^2 − 6y = −4 Then separate them into groups (x^2 + 4x) + (y^2-6y) = -4 Making them trinomials and squaring (x^2+4x+4) + (y^2 - 6y+9) = -4 + 4+9 Then we can change the trinomials, factor, and simplify the right (x+2)^2 + (y-3)^2 = 9 So it converts to (x-h)^2 + (y-k)^2 =r^2 (x+2)^2 + (y-3)^2 = 3^2 Where (h,k) is the center, r is radius So then we have (-2,3) is center, and 3 is radius

snowflake0531:

Change Making them trinomials and squaring to making them trinomials so we are able to factor it also change So it converts to (x-h)^2 + (y-k)^2 =r^2 (x+2)^2 + (y-3)^2 = 3^2 it's not converting We have (x-h)^2 + (y-k)^2 = r^2 as the base form, so with this part: Where (h,k) is the center, r is radius" and "(x+2)^2 + (y-3)^2 = 3^2" this part we can determine the center point and radius

Seafoam:

Ok I changed those two sentences

snowflake0531:

Kkay, then the explanation stuff is great xd

Seafoam:

Whoo thank god

Seafoam:

On to the next one 😅

snowflake0531:

xdddd k

Seafoam:

Just 4 more

Seafoam:

And then another assignment..

snowflake0531:

Kay I'll help one what i can xd